OpenStudy (anonymous):

determine whether the pair of line is parallel, perpendicular ,or neither x-6y=-8 y=3x-6

ganeshie8 (ganeshie8):

rewrite the given lines in below form : \(y = mx + b\)

ganeshie8 (ganeshie8):

\(m\) is the slope

ganeshie8 (ganeshie8):

if slopes are \(equal\), then lines are \(parallel\) if product of slopes is \(-1\), then lines are \(perpendicular\) if it doesnt belong to any one of above two, then its \(neither\)

OpenStudy (*insert name here*):

^Answer right there.. what more do you need?

OpenStudy (anonymous):

i don' understand,and i have to show all working.

OpenStudy (nincompoop):

do you know how to add, subtract, multiply and divide?

OpenStudy (anonymous):

yes

OpenStudy (nincompoop):

solve for y x-6y=-8

OpenStudy (anonymous):

y=1/6

OpenStudy (nincompoop):

what happened to your x?

OpenStudy (nincompoop):

and I thought you said you knew division :(

OpenStudy (anonymous):

in my defense i did say i don't understand!

OpenStudy (nincompoop):

do it again, solve for y x-6y=-8